### v1

@@ 3.4 Question Generation
Each cleaned and preprocessed content item—whether a word, phrase, or proverb—was programmatically paired with its dialect label, meaning, and usage example to construct a complete input prompt. This setup enabled GPT-4o to generate questions that evaluated not only surface-level understanding but also deeper comprehension of dialectal and cultural nuances. For each dialectal item, the model generated six distinct types of questions, with each type targeting a specific dimension of language understanding:
@@
This systematic generation approach ensured comprehensive coverage of the linguistic and cultural features embedded in Saudi dialects. By applying the process across all content categories—words, phrases, and proverbs—a total of 18,564 multiple-choice questions were generated. These questions form the core of the Absher benchmark and enable a rigorous, multifaceted evaluation of LLM performance. Table 3 shows the distribution of questions across regional dialects, highlighting the benchmark’s broad linguistic representation.
@@
Figure 4 illustrates the six question types included in the benchmark. Together, these types enable a well-rounded evaluation of model capabilities, covering meaning interpretation, cultural and regional understanding, and recognition of local expressions. The use of a standardized prompt template throughout the generation process ensures consistency, minimizes bias, and strengthens the reliability of the evaluation.
@@ 3.5 Human Evaluation
To validate the reliability, correctness, and cultural appropriateness of the questions in the Absher benchmark, we conducted a structured human evaluation. This step was essential to ensure that the automatically generated content adhered to linguistic standards and dialectal relevance, especially given the complexity and regional variability of Saudi dialects. Four native Saudi Arabic speakers participated in the evaluation. All annotators were fluent in multiple regional dialects and possessed strong familiarity with the cultural and linguistic diversity across Saudi Arabia. They were divided into two independent groups, each assigned a distinct subset of the data. This setup enabled broader dataset coverage, minimized cross-group bias, and reduced cognitive fatigue.
A pilot study preceded the main annotation phase to ensure consistent interpretation of the evaluation guidelines. Annotators were presented with 20 representative sample contents (120 questions total), covering diverse dialects and question types. These examples served to familiarize them with the annotation tasks and resolve any ambiguities in advance. To test attentiveness and accuracy, the pilot set included quality control questions with plausible yet incorrect answers. Annotators who misclassified these were given follow-up clarifications to align their understanding with the evaluation standards. This step significantly improved consistency and judgment reliability in the formal evaluation.
The main evaluation involved a stratified random sample comprising 10% of the benchmark contents—Words, Phrases, and Proverbs—across all six question types. This resulted in 600 annotated questions. Complete annotation guidelines, including task definitions, assessment criteria, and illustrative examples, are provided in Appendix B. Based on these instructions, annotators completed two core tasks:
@@
Evaluation Meaning True/False FillB Cont.Usage Cult.Interpretation Loca. Recognition
Q-Quality 0.97 0.95 0.71 0.72 0.83 0.85
A-Correctness 0.97 0.87 0.91 0.88 0.94 0.92
Table 4: Cohen’s Kappa scores across the two evaluated criteria. FillB, Cont., Cult., and Loca. refer to “contextual”, “cultural”, “location”, and “Fill-in-the-Blank”.
@@ 4 Experiments
We evaluated a set of open-source LLMs: LLaMA-3 8B Instruct [25], Jais-13B [26], ALLaM-7B [27], Mistral-7B [28], Qwen2.5-7B Instruct [29], and AceGPT-7B-chat [30]. These models were chosen based on three main criteria: open availability, Arabic language support, and competitive performance across a range of NLP tasks. All models were evaluated in a zero-shot setting, without any task-specific fine-tuning, to assess their inherent ability to understand Saudi dialects and cultural content. Since the benchmark is composed exclusively of multiple-choice questions, we employed four standard classification metrics: Accuracy, Precision, Recall, and F1 Score. These metrics offer a comprehensive and balanced evaluation of model performance, in line with prior work on zero-shot LLMs assessment for low-resource languages [31].
@@
Table 5: Average Evaluation Results of Models Performance
@@ 4.1 Results
Table 5 presents the performance of the six evaluated models—ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT—across different content types, namely words, phrases, and proverbs. The results reveal noticeable variations in performance depending on both the linguistic content and the complexity of the task [^2] The full set of results is provided in Appendix C..
For word-level content, the Qwen model demonstrated the strongest capabilities, achieving an accuracy of 63.00% and an F1-score of 41.18%. These results indicate Qwen’s effectiveness in handling isolated lexical items, suggesting its advantage in processing tasks that rely on clear, context-independent meanings. In contrast, phrase-level tasks presented a more competitive landscape. Mistral achieved the highest F1-score at 26.02%, with Qwen and ACEGPT following closely behind. This reflects a broader challenge for all models when dealing with phrases, which often carry more ambiguous or context-dependent meanings compared to individual words. Proverb-based tasks emerged as the most challenging across all models, reflecting the complexity and cultural depth embedded in such expressions. Despite these challenges, ALLaM led in this category with an accuracy of 48.74% and an F1-score of 31.43%, indicating its relative strength in handling idiomatic and culturally grounded language. In contrast, models such as Jais consistently recorded lower F1-scores across all content types, particularly for proverbs, highlighting limitations in their ability to process culturally rich or metaphorical expressions.
@@ 5.1 Impact of Model Size and Language Orientation
Model Size Language Accuracy [Uncaptioned image] Qwen2.5 Instruct 1st 7B Multilingual 50.35% [Uncaptioned image] Mistral 2nd 7B Multilingual 45.93% [Uncaptioned image] ALLaM 3rd 7B Arabic 43.78% [Uncaptioned image] LLaMA-3 Instruct 8B Multilingual 43.34% [Uncaptioned image] Jais 13B Arabic 42.14% [Uncaptioned image] AceGPT-chat 7B Arabic 38.88%
Table 6: Models Overview, and Average accuracy
@@ 5.2 Examination of Model Responses by Region
Our evaluation of Large Language Models (LLMs) responses to region-specific linguistic questions, detailed in Figure 5, reveals considerable performance variability influenced by both the geographic origin of expressions and the nature of the linguistic task. Notably, we observed strong performance in the General category for a widely known proverb, with five out of six models providing correct answers. This aligns with the expectation that frequently cited expressions, which appear widely in public discourse, are more likely to be accurately interpreted due to their higher representation in pretraining datasets.
@@ 5.3 Performance Across Content and Question Types
5.4 Case Study of LLMs Responses Across Question Types and Dialects
Our case study investigates the performance of six LLMs across culturally embedded tasks derived from various Saudi dialectal regions. To illustrate observed performance patterns, we examine specific model responses across varied content types, question structures, and dialects, as detailed in Tables 7 to 9. The results indicate that regionally common and widely used expressions are more accurately interpreted by LLMs. For example, in the general category, the proverb:
{RLtext}
“إذا حجت البقرة على قرونها”
“If the cow performs Hajj on its horns” (Table 7) was correctly understood by five out of six models: ALLaM, LLaMA, Jais, Qwen, and ACEGPT all provided the correct answer, while only Mistral failed. This proverb is relatively widespread, and its meaning—“something impossible to achieve”—is often documented in public sources, likely increasing its representation in model training data. In contrast, proverbs that are more regionally isolated or culturally specific, such as the Southern fill-in-the-blank proverb :
{RLtext}
“خطط في ماء واقبص في حيد”
“Plan in water and pinch in a rock” (Table 8), saw a complete failure across all six models. None of the models—ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT selected the correct option, suggesting a consistent gap in exposure to underrepresented regional idioms, especially those originating from areas with fewer accessible textual resources.
Additionally, the study revealed notable differences in the models’ ability to comprehend regional dialects, particularly in how they responded to specific tasks based on region and question type. For example, when tasked with identifying a location-specific term (Table 9) from the Central region, only two models, ALLaM and Jais, were able to correctly associate it with the Al-Qassim region. The other models, however, defaulted to more familiar urban areas, like Riyadh and Jeddah, likely due to the prevalence of these dialects in the training materials. Urban dialects from cities such as Riyadh, Jeddah, and Khobar are more commonly represented in online resources, which could explain why the models showed a preference for these regions. In contrast, like Qassim, Aseer, or the southern regions face challenges due to limited textual data from these areas.
@@ 6 Discussion
This research presents Absher, a novel benchmark aimed at evaluating LLMs on Saudi dialects—a linguistically rich yet underrepresented subset of Arabic. Unlike prior work such as Dallah [14], Palm [18], and AL-QASIDA [22], which aggregate dialects or focus on broader Arabic settings, Absher uniquely provides fine-grained, region-specific coverage of Saudi dialects, including Central, Weastern, Southern, Eastern, and Northern varieties.
@@
Our findings highlight notable patterns in how LLMs handle Saudi dialects. Location Recognition tasks yielded the highest performance, likely due to the concrete and frequently occurring nature of geographic references in training data, which aids retrieval and identification. In contrast, True/False questions were the most challenging—especially in low-context settings—reflecting the inherent difficulty of binary reasoning tasks that require nuanced cultural and contextual understanding beyond surface-level lexical cues.
@@
Furthermore, Qwen’s strong performance—especially in Meaning and Contextual Usage questions—demonstrates its robust semantic understanding and ability to accurately interpret dialectal expressions in context. Across content types, tasks involving Proverbs consistently achieved higher accuracy than those based on individual Words or Phrases. This suggests that models benefit from the richer contextual and cultural cues provided by longer, more nuanced inputs. For instance, ALLAM’s relatively high performance in Cultural Interpretation of proverbs highlights its capacity for deeper socio-cultural reasoning, likely attributed to its Arabic-centric training.
@@
The results reveal a noticeable gap between multilingual and Arabic-native models. Models like Qwen and Mistral, which were trained on a wide range of languages, outperformed their Arabic-centric counterparts in overall accuracy. Their ability to generalize across dialectal variations, despite not being trained explicitly on Saudi dialects, reflects the benefits of exposure to diverse language data. These results align with observations from Palm [18] and related studies, where multilingual training correlated with higher robustness on non-standard Arabic input.
@@
Nevertheless, Arabic-native models demonstrated strengths in specific areas. For example, ALLaM achieved high accuracy in tasks involving proverbs, while Jais performed relatively well in location-based recognition. These patterns suggest that domain-specific pretraining provides advantages in interpreting traditional or idiomatic expressions, even if it does not generalize as effectively across diverse question types. Such findings align with results from the AraDiCE benchmark [3], where models exhibited improved performance when test data resembled the training distribution.
To further demonstrate model differences, consider a Cultural Interpretation question based on the proverb:
{RLtext}
“اللي ما يعرف الصقر يشويه”
“He who doesn’t recognize the falcon cooks it”
While ALLAM accurately selected the metaphorical interpretation (“Those who lack knowledge may misuse valuable things”), other models like Jais and ACEGPT selected overly literal or generic options. This reflects ALLAM’s stronger grounding in socio-cultural expressions—a critical strength when applying LLMs in culturally sensitive settings.
@@
In summary, our findings reveal a clear gap in current LLMs’ ability to comprehensively handle the diverse spectrum of Saudi Arabic dialects, especially in less standardized or culturally nuanced forms. While multilingual models demonstrate stronger generalization capabilities, Arabic-native models show targeted strengths in specific areas such as cultural interpretation. However, neither category achieves consistently strong performance across all dialects and task types. These insights underscore the need for more representative and balanced datasets, as well as tailored adaptation strategies, to improve model alignment with dialectal Arabic at the national level.
7 Ethical Considerations and Limitations

### v2

@@ 3.4 Question Generation
Each cleaned and pre-processed content item, whether a word, phrase, or proverb, was programmatically paired with its dialect label, meaning, and usage example to construct a complete input prompt. This setup enabled GPT-4o to generate questions that evaluated not only surface-level understanding but also deeper comprehension of dialectal and cultural nuances. For each dialectal item, the model generated six distinct types of questions, with each type targeting a specific dimension of language understanding:
@@
This systematic generation approach ensured comprehensive coverage of the linguistic and cultural characteristics embedded in the Saudi dialects. By applying the process across all content categories, words, phrases, and proverbs, a total of 18,564 multiple-choice questions were generated. These questions form the core of the Absher benchmark and allow for a rigorous, multifaceted evaluation of LLM performance. Table 3 shows the distribution of questions across regional dialects, highlighting the benchmark’s broad linguistic representation.
The distribution of questions across regional dialects reflects the linguistic landscape of Saudi Arabia. The Najdi dialect, spoken predominantly in the Central region, and the Hejazi dialect, prevalent in the Western region, are spoken by a larger portion of the population and receive extensive exposure through media and online platforms. This widespread usage provides a richer pool of linguistic material. By contrast, the Southern, Northern, and Eastern dialects have a more localized presence, with fewer readily accessible resources. Despite these differences, the benchmark includes all major dialects, thus capturing the full spectrum of Saudi Arabic and supporting a comprehensive evaluation of LLM performance.
@@
Figure 3 illustrates the six types of questions included in the benchmark. Together, these types enable a well-rounded evaluation of model capabilities, covering meaning interpretation, cultural and regional understanding, and recognition of local expressions. The use of a standardized prompt template throughout the generation process ensures consistency, minimizes bias, and strengthens the reliability of the evaluation.
@@ 3.5 Human Evaluation
To validate the reliability, correctness, and cultural appropriateness of the questions in the Absher benchmark, we performed a structured human evaluation. This step was essential to ensure that the automatically generated content adhered to linguistic standards and dialectal relevance, especially given the complexity and regional variability of Saudi dialects. Four native Saudi Arabic speakers participated in the evaluation. All annotators were fluent in multiple regional dialects and possessed strong familiarity with the cultural and linguistic diversity across Saudi Arabia. They were divided into two independent groups, each assigned a distinct subset of the data. This setup enabled broader dataset coverage, minimized cross-group bias, and reduced cognitive fatigue.
A pilot study preceded the main annotation phase to ensure consistent interpretation of the evaluation guidelines. Annotators were presented with 20 representative sample contents (120 questions total), covering diverse dialects and question types. These examples served to familiarize them with the annotation tasks and resolve any ambiguities in advance. To test attentiveness and accuracy, the pilot set included quality control questions with plausible yet incorrect answers. Annotators, who misclassified these, were received follow-up clarifications to align their understanding with the evaluation standards. This step improved consistency and reliability of the judgment in the formal evaluation.
The main evaluation involved a stratified random sample comprising 10% of the benchmark contents, i.e., Words, Phrases, and Proverbs, across all six question types. This resulted in 600 annotated questions. Complete annotation guidelines, including task definitions, assessment criteria, and illustrative examples, are provided in Appendix B. Based on these instructions, annotators completed two core tasks:
@@
Table 4: Cohen’s Kappa scores for question quality and answer correctness across different question types. Bold denotes the top scores.
Q-Type Q-Quality A-Correctness
Meaning 0.97 0.97
True/False 0.95 0.87
Fill-in-the-Blank 0.71 0.91
Contextual Usage 0.72 0.88
Cultural Interpretation 0.83 0.94
Location Recognition 0.85 0.92
@@ 4 Experiments
We evaluated a set of open-source LLMs: LLaMA-3 8B Instruct [Touvron2024], Jais-13B [sengupta2023], ALLaM-7B [bari2025allam], Mistral-7B [jiang2023], Qwen2.5-7B Instruct [yang2024qwen2p5], and AceGPT-7B-chat [huang2023acegpt]. These models were chosen based on three main criteria: open availability, Arabic language support, and competitive performance across a range of NLP tasks. All models were evaluated in a zero-shot setting, without any task-specific fine-tuning, to assess their inherent ability to understand Saudi dialects and cultural content. Since the benchmark is composed exclusively of multiple-choice questions, we employed four standard classification metrics: Accuracy, Precision, Recall, and F1 Score. These metrics offer a comprehensive and balanced evaluation of model performance, in line with prior work on zero-shot LLMs assessment for low-resource languages [lowres_llm_eval_2025].
Table 5: Average evaluation results of models performance. Bold denotes the best result achieved across all models for each content type, while underlining indicates the best-performing content type for each individual model.
@@ 4.1 Results
Table 5 presents the performance of the six evaluated models, such as ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT, across different content types, namely words, phrases, and proverbs. The results reveal noticeable variations in performance depending on both the linguistic content and the complexity of the task [^2] The full set of results is provided in Appendix D.. For word-level content, the Qwen model demonstrated the strongest capabilities, achieving an accuracy of 63.00% and an F1-score of 41.18%. These results indicate Qwen’s effectiveness in handling isolated lexical items, suggesting its advantage in processing tasks that rely on clear, context-independent meanings.
To contextualize the reported performance, it is important to note the chance-level baseline for this benchmark. Since most multiple-choice questions contain four options, random guessing corresponds to 25% accuracy (and 50% for True/False items). Therefore, the observed performance range of approximately 40–60% accuracy across models is substantially above chance, highlighting both the non-trivial capability of the evaluated LLMs and the inherent difficulty of the dialect- and culture-sensitive questions in Absher.
In contrast, phrase-level tasks presented a more competitive landscape. Mistral achieved the highest F1-score at 26.02%, with Qwen and ACEGPT following closely behind. This reflects a broader challenge for all models when dealing with phrases, which often carry more ambiguous or context-dependent meanings compared to individual words. Proverb-based tasks emerged as the most challenging across all models, reflecting the complexity and cultural depth embedded in such expressions. Despite these challenges, ALLaM led in this category with an accuracy of 48.74% and an F1-score of 31.43%, indicating its relative strength in handling idiomatic and culturally grounded language. In contrast, models such as Jais consistently recorded lower F1-scores across all content types, particularly for proverbs, highlighting limitations in their ability to process culturally rich or metaphorical expressions.
@@ 5.1 Impact of Model Size and Language Orientation
Table 6: Overview of the evaluated models, highlighting their size, supported languages, and average accuracy achieved in the experiments. The best results are listed from highest to lowest, with the three best results in bold.
Model Size Language Accuracy
[Uncaptioned image] Qwen2.5 Instruct 7B Multilingual 50.35%
[Uncaptioned image] Mistral 7B Multilingual 45.93%
[Uncaptioned image] ALLaM 3rd 7B Arabic 43.78%
[Uncaptioned image] LLaMA-3 Instruct 8B Multilingual 43.34%
[Uncaptioned image] Jais 13B Arabic 42.14%
[Uncaptioned image] AceGPT-chat 7B Arabic 38.88%
@@ 5.2 Examination of Model Responses by Region
Our evaluation of LLMs responses to region-specific linguistic questions, detailed in Figure 4, reveals considerable performance variability influenced by both the geographic origin of expressions and the nature of the linguistic task. Notably, we observed strong performance in the General category for a widely known proverb, with five out of six models providing correct answers. This aligns with the expectation that frequently cited expressions, which appear widely in public discourse, are more likely to be accurately interpreted due to their higher representation in pretraining datasets.
@@ 5.3 Performance Across Content and Question Types
5.4 Case Study of LLMs Responses Across Question Types and Dialects
Our case study investigates the performance of six LLMs across culturally embedded tasks derived from various Saudi dialectal regions. To illustrate observed performance patterns, we examine specific model responses across varied content types, question structures, and dialects, as detailed in Tables 7 to 9. The results indicate that regionally common and widely used expressions are more accurately interpreted by LLMs. For example, in the general category, the proverb: ”Itha Hajat Al-Baqara Ala Qoroniha”, its translation is ”If the cow performs Hajj on its horns” (See Table 7) was correctly understood by five out of six models: ALLaM, LLaMA, Jais, Qwen, and ACEGPT all provided the correct answer, while only Mistral failed. This proverb is relatively widespread, and its meaning, “something impossible to achieve”, is often documented in public sources, likely increasing its representation in model training data. In contrast, proverbs that are more regionally isolated or culturally specific, such as the Southern fill-in-the-blank proverb : ”Khattet fi maa wa aqbas fi heid”, its translation is ”Plan in water and pinch in a rock” (See Table 8), saw a complete failure across all six models. None of the models, i.e., ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT, selected the correct option, suggesting a consistent gap in exposure to underrepresented regional idioms, especially those originating from areas with fewer accessible textual resources.
@@ 5.4 Case Study of LLMs Responses Across Question Types and Dialects
Additionally, the study revealed notable differences in the models’ ability to comprehend regional dialects, particularly in how they responded to specific tasks based on region and question type. For example, when tasked with identifying a location-specific term (See Table 9) from the Central region, only two models, ALLaM and Jais, were able to correctly associate it with the Al-Qassim region. The other models, however, defaulted to more familiar urban areas, like Riyadh and Jeddah, likely due to the prevalence of these dialects in the training materials. Urban dialects from cities such as Riyadh, Jeddah, and Khobar are more commonly represented in online resources, which could explain why the models showed a preference for these regions. In contrast, like Qassim, Aseer, or the southern regions face challenges due to limited textual data from these areas.
@@ 6 Discussion
This research presents Absher, a novel benchmark aimed at evaluating LLMs on Saudi dialects, a linguistically rich yet underrepresented subset of Arabic. Unlike prior work such as Dallah [alwajih2024dallah], Palm [alwajih2025palm], and AL-QASIDA [robinson2024qasida], which aggregate dialects or focus on broader Arabic settings, Absher uniquely provides fine-grained, region-specific coverage of Saudi dialects, including Central, Western, Southern, Eastern, and Northern varieties. Absher’s strength lies in its foundation: over 18,000 multiple-choice questions derived from a curated dataset of dialectal words, phrases, and proverbs sourced from these five major Saudi regions. The benchmark systematically spans six distinct task categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
@@
Our findings highlight notable patterns in how LLMs handle Saudi dialects. Location Recognition tasks yielded the highest performance, likely due to the concrete and frequently occurring nature of geographic references in training data, which aids retrieval and identification. In contrast, True/False questions were the most challenging, especially in low-context settings, reflecting the inherent difficulty of binary reasoning tasks that require nuanced cultural and contextual understanding beyond surface-level lexical cues.
@@
Furthermore, Qwen’s strong performance, especially in Meaning, Contextual Usage, and Cultural Interpretation questions, demonstrates its robust semantic understanding and ability to accurately interpret dialectal expressions in context. Across content types, performance varied by model rather than being uniformly highest for any single type. While some models (e.g., ALLaM) achieved relatively stronger results on Proverbs, others, such as Qwen, performed best on Word-level questions, indicating that certain models benefit more from the richer contextual and cultural cues provided by longer and more nuanced inputs. For instance, ALLAM’s relatively high performance in Cultural Interpretation of proverbs highlights its ability for deeper socio-cultural reasoning, likely attributed to its Arabic-centric training.
@@
The results reveal a noticeable gap between multilingual and Arabic-native models. Models like Qwen and Mistral, which were trained on a wide range of languages, outperformed their Arabic-centric counterparts in overall accuracy. Their ability to generalize across dialectal variations, despite not being trained explicitly on Saudi dialects, reflects the benefits of exposure to diverse language data. These results align with observations from Palm [alwajih2025palm] and related studies, where multilingual models correlated with higher robustness on non-standard Arabic input.
@@
Nevertheless, Arabic-native models demonstrated strengths in specific areas. For example, ALLaM achieved high accuracy in tasks involving proverbs, while Jais performed relatively well in location-based recognition. These patterns suggest that domain-specific pretraining provides advantages in interpreting traditional or idiomatic expressions, even if it does not generalize as effectively across diverse question types. Such findings align with results from the AraDiCE benchmark [Baly2023], where models exhibited improved performance when test data resembled the training distribution. To further demonstrate model differences, consider a Cultural Interpretation question based on the proverb: ”Ally ma ya’arf alsger yshwaeh”, meaning ”He who doesn’t recognize the falcon cooks it” While ALLAM accurately selected the metaphorical interpretation (“Those who lack knowledge may misuse valuable things”), other models like Jais and ACEGPT selected overly literal or generic options. This reflects ALLAM’s stronger grounding in socio-cultural expressions—a critical strength when applying LLMs in culturally sensitive settings.
@@
In summary, our findings reveal a clear gap in current LLMs’ ability to comprehensively handle the diverse spectrum of Saudi Arabic dialects, especially in less standardized or culturally nuanced forms. While multilingual models demonstrate stronger generalization capabilities, Arabic-native models show targeted strengths in specific areas such as cultural interpretation. However, neither category achieves consistently strong performance across all dialects and task types. This discrepancy can be attributed to structural imbalances in the representation of dialectal forms within the training data, where widely documented or urban-centered dialects are more prevalent, whereas regionally specific or culturally embedded expressions remain underrepresented. Furthermore, performance patterns indicate that model success is influenced not only by linguistic coverage but also by the alignment between task type and the type of cultural knowledge required. These insights underscore the need for more representative and balanced datasets, alongside tailored adaptation strategies and controlled evaluation methods, to improve model alignment with dialectal Arabic at the national level and to better understand why certain models excel in some contexts while struggling in others.
6.1 Limitations
While the Absher benchmark represents a significant step toward evaluating LLMs on Saudi dialects and culture, the study is subject to several limitations. First, the evaluations were conducted under constrained computational resources, which limited our ability to explore multiple model configurations or conduct repeated trials. Second, all models were assessed using a fixed prompt structure (e.g., “Answer as if you are a Saudi…”), which may have influenced model behavior. Future work could explore the impact of prompt variation and instruction tuning. Third, although human evaluation was performed on a stratified sample to ensure dataset quality, a complete manual review of all 18,564 questions was not feasible due to the scale and diversity of the benchmark. Finally, while Absher covers various regional Saudi dialects, the representation of these dialects within the dataset is not perfectly balanced; for instance, Central and Western dialects are more prevalent. That highlights the need for more diverse and evenly distributed resources to comprehensively capture all Saudi dialectal variations in future research. Future research should leverage more extensive computational infrastructure, experiment with prompt engineering strategies and fine-tuning methods. It should also incorporate broader expert reviews to further enhance the accuracy, reliability, and cultural validity of the benchmark evaluations.